OpenStudy (anonymous):

36% of what number is 180?

OpenStudy (shaik0124):

64.8 is answer

OpenStudy (anonymous):

Let the number be: x : \(36 \% \; \; of \; \; x = 180\) Or : \[\frac{36}{100} \times x = 180\] Can you find x from here...

OpenStudy (anonymous):

\[\frac{ 36 *x}{ 100 } = 180\] find x

OpenStudy (anonymous):

what does 100 have to do with it?

OpenStudy (anonymous):

It will be better if you will give solution along with your direct answer.. @shaik0124

OpenStudy (anonymous):

% means 100..

OpenStudy (anonymous):

If I say 45% of anything,then we assume that whole thing as 100 and out of which 45 is the fraction...

OpenStudy (anonymous):

all i know is 64.8 isnt the answer

OpenStudy (anonymous):

Say, if I say 10% of 100.. Then it simply means, that out of 100, 10% we have taken. And 10 % of 100 is 10 only... So, 10 out of 100 it basically means.

OpenStudy (anonymous):

i found it thanks for your help

OpenStudy (anonymous):

Please let me know if you are not getting... I will explain it in different way then... And you are welcome dear..

OpenStudy (shaik0124):

nice punjabi :)

OpenStudy (anonymous):

From INDIA??? @shaik0124

OpenStudy (shaik0124):

ha ha

OpenStudy (shaik0124):

where do u study

OpenStudy (anonymous):

At openstudy.com Ha ha ha...